# ReinWiFi: A Reinforcement-Learning-Based Framework for the Application-Layer QoS Optimization of WiFi Networks

### v1

@@ Abstract
In this paper, a reinforcement-learning-based scheduling framework is proposed and implemented to optimize the application-layer quality-of-service (QoS) of a practical wireless local area network (WLAN) suffering from unknown interference. Particularly, application-layer tasks of file delivery and delay-sensitive communication, e.g., screen projection, in a WLAN with enhanced distributed channel access (EDCA) mechanism, are jointly scheduled by adjusting the contention window sizes and application-layer throughput limitation, such that their QoS, including the throughput of file delivery and the round trip time of the delay-sensitive communication, can be optimized. Due to the unknown interference and vendor-dependent implementation of the network interface card, the relation between the scheduling policy and the system QoS is unknown. Hence, a reinforcement learning method is proposed, in which a novel Q-network is trained to map from the historical scheduling parameters and QoS observations to the current scheduling action. It is demonstrated on a testbed that the proposed framework can achieve a significantly better QoS than the conventional EDCA mechanism.
@@ I Introduction
Reinforcement learning (RL) for radio resource management has been attracting tremendous attention since it is a promising technique to tackle unknown system statistics and solve the prohibitive policy optimization problem with tolerable complexity and good performance. Moreover, the RL technique also has great potential to optimize a wireless system even without accurate or complete observation of the system state, which might happen in practical implementations.
@@
There have been a significant amount of works optimizing the throughput, delay or age-of-information (AoI) of wireless networks via the method of RL. Most of these works assumed full knowledge of the system state in algorithm design, which could be applied to the systems where the global system state could be collected at a centralized controller. On the other hand, RL was also utilized to optimize the performance of wireless systems with distributive transmission scheduling, e.g., wireless fidelity (WiFi) systems. For instance, an adaptive channel contention mechanism was proposed for WiFi systems in [1], where a local RL agent was deployed at each user equipment (UE). The local agents adjusted the minimum contention window (MCW) size according to the global statistics of successful channel contention such that the transmission fairness among the agents can be ensured. Instead of global statistics, a distributive RL algorithm with the assistance of federated learning was proposed in [2] to adapt the channel contention according to the local channel state, such that the local throughput was optimized. Moreover, a deep multi-agent RL technique based on the QMIX algorithm [3] was proposed in [4] to improve network throughput while maintaining user fairness. In this work, the channel contention decision was made according to the history of the last transmission duration. In order to resolve the collision issue of the distributive channel access, deep RL algorithms were proposed in [5] to determine the timing of doubling the contention window based on the estimated collision probability. In addition to the adaptive channel contention, a double deep Q-network (DDQN) [6] based rate adaptation algorithm was proposed in [7] to improve network throughput, where the agent learned the optimal transmission rate based on the modulation and coding scheme (MCS) and frame loss rate. Most of the above literature assumed knowledge of the physical (PHY) layer and media access control (MAC) layer system states. However, it might be challenging to obtain such knowledge in the scheduler design of a practical WiFi network. Moreover, the absence of knowledge on co-channel interference and the vendor-dependent implementation of WiFi adapters would also raise challenges in the optimization of scheduling policies.
@@
In this paper, we would like to shed some light on the RL-based scheduling design for practical WiFi systems suffering from unknown co-channel interference. Particularly, a framework, namely ReinWiFi, is proposed for the scheduling of delay-sensitive communication tasks and file delivery tasks in the application layer of a WiFi network. In ReinWiFi, a controller periodically collects the past scheduling parameters and average quality-of-service (QoS) observations of all the application-layer tasks, determines rate limitation and contention window size for all the transmitters, such that the total throughput of file delivery tasks is maximized and the latency requirements of delay-sensitive tasks are ensured. It is shown by the experiments that the proposed framework can adapt to the variation of task number, interfering traffic, and link quality, and significantly outperforms the conventional EDCA mechanism.
@@ II-A Deployment Scenario
Because of the transmission latency constraint, the delay-sensitive tasks should be scheduled with higher priority than the file delivery ones. Hence, all the transmitters access the channel via the enhanced distributed channel access (EDCA) mechanism defined in IEEE 802.11e. Particularly, four access category (AC) queues, namely voice (VI), video (VO), best effort (BE), and background (BK), are adopted at all the transmitters. The transmission priorities of the four AC queues are differentiated by values of arbitration inter-frame spacing (AIFS) and contention window (CW) size. As in the practical systems, the file delivery tasks are scheduled with the BE priority, and the delay-sensitive tasks are scheduled with the VI priority. The latter has smaller AIFS and CW size, leading to a larger successful probability in channel contention. As a remark, due to the distributive channel contention mechanism, it is infeasible to accurately control the packet transmission order among the devices of a WiFi network with commercial network interface cards (NICs). Instead, the packet transmission in the ReinWiFi system is scheduled in a stochastic manner by adapting the CW sizes of AC queues in each device.
@@ II-B Task Queuing Model
For each file delivery task, all the information bits to be delivered are saved in an application-layer buffer, and a user datagram protocol (UDP) socket is established at the very beginning of transmission. The data dispatch from the buffer to the UDP socket is controlled by a dispatcher. The UDP socket encapsulates the received data from the dispatcher into UDP datagrams and forwards them to the driver of NIC for WiFi transmission accordingly. As a remark, the new datagrams at the NIC may not be transmitted immediately. In fact, each NIC maintains four MAC-layer AC queues associated with the four transmission priorities, respectively. The arrival datagrams are saved in the corresponding queues and transmitted following the vendor’s protocol. The queuing status of the NIC is usually not accessible in the application-layer. Thus, it is infeasible for the proposed system to know when the NIC completely delivers a datagram; it is, therefore, infeasible for the proposed system to precisely control the transmission of a UDP datagram or an application-layer packet. As a result, the scheduling of the proposed system is designed based on the average observable performance in the application layer.
Specifically, the transmission time is organized into a sequence of scheduling periods, each with a duration of $T_{s}$ seconds. $T_{s}$ is sufficiently large to accommodate a number of MAC protocol data unit transmissions. Due to the invisibility of NIC status, the QoS of a file delivery task is measured by its application-layer throughput in one scheduling period. Particularly, for the ${m}$-th file delivery task of the $({i},{j})$-th link, its QoS in the $t$-th scheduling period $r_{{i},{j}}^{m}(t)$ is defined as the number of information bits transferred from the task buffer to the associated UDP socket. The dispatcher is designed to adaptively limit the throughput of the file delivery task such that delay-sensitive tasks could have a larger chance to access the channel. Hence, let $b_{{i},{j}}^{m}(t)$ be the throughput limitation of the ${m}$-th file delivery task of $({i},{j})$-th link in the $t$-th scheduling slot, the dispatcher would make sure
@@
For each delay-sensitive task, a task queue and UDP socket are established at the very beginning. The application-layer packets arrive at the task queue periodically with a fixed average data rate. The first packet in the queue is forwarded to the UDP socket for WiFi transmission as long as the socket is idle. Due to the lack of MAC-layer status, the measurement of the transmission latency of a packet could hardly be accurate. Hence, we use the round-trip time (RTT) as the QoS measurement of delay-sensitive communication tasks. Particularly, for each delay-sensitive task, an acknowledgment will be sent back from the receiver to the transmitter when an application-layer packet is completely received. Hence, the transmitter can calculate the RTTs of all packet transmissions. For the ${m}$-th delay-sensitive communication task of the $({i},{j})$-th link ($\forall({i},{j})\in{\mathcal{L}},{m}\in{\mathcal{T}_{{i},{j}}^{r}},$), its QoS in the $t$-th scheduling period $d_{{i},{j}}^{m}(t)$ is defined as the average RTT of the packets transmitted in this scheduling period.
@@ II-C Scheduling Model
Denote the CW sizes of the VI and BE priorities of the ${i}$-th device at $t$-th scheduling period as $w_{i}^{\mathtt{VI}}(t)$ and $w_{i}^{\mathtt{BE}}(t)$ respectively, we shall focus on the joint scheduling of these channel contention parameters as well as the dispatchers’ throughput limitation $\{b_{{i},{j}}^{m}(t)|\forall({i},{j})\in{\mathcal{L}},{m}\in{\mathcal{T}_{{i},{j}}^{f}}\}$ in each scheduling period.
Particularly, each transmitter collects the QoS observations of its tasks at the end of each scheduling period and delivers them to a centralized controller, which can be implemented in an AP or other device. Not all the tasks in the universal task sets are in the active state. The average RTTs and throughputs of the inactive delay-sensitive and file delivery tasks are represented by a sufficiently large value and 00, respectively. Hence, the aggregation of QoS observations received at the controller at the end of the $t$-th scheduling period can be represented as
@@
Due to the time-varying traffic of the interfering devices, the scheduling parameters, including the file throughput limitations and CW sizes, are adapted at the centralized controller in each schedule according to the system’s scheduling parameters and QoS observations in the past $N$ scheduling periods. Specifically, the aggregation of scheduling parameters over a period is represented as
@@
Thus, at the very beginning of the $t$-th scheduling period, $\mathcal{A}_{t}$ ($\forall t$) is determined based on past scheduling parameters and QoS observations $\{(\mathcal{O}_{t-N},\mathcal{A}_{t-N}),(\mathcal{O}_{t-N+1},\mathcal{A}_{t-N+1}),\ldots,(\mathcal{O}_{t-1},\mathcal{A}_{t-1})\}$.
@@ III Problem Formulation
The proposed ReinWiFi system should successively make scheduling decisions for each scheduling period. Hence, it is formulated as a Markov decision process (MDP) in the following.
@@
Denote $\mathcal{A}_{t}$ defined in (3) as the scheduling action in the $t$-th scheduling period, $\mathcal{A}_{t}^{i}\triangleq\left\{b_{i},{j}^{m}(t)|\forall j\in\mathcal{L}_{i},{m}\in{\mathcal{T}_{{i},{j}}^{f}}\right\}\cup\left\{w_{i}^{\mathtt{VI}}(t),w_{i}^{\mathtt{BE}}(t)\right\}$, as the local scheduling action of the $i$-th device in the $t$-th scheduling period. The scheduling policy $\Omega$ is a mapping from state space to action space as $\Omega(\mathcal{S}_{t})=\mathcal{A}_{t}$.
@@
where $\omega$ is a weight, $\mathrm{D}_{{i},{j}}^{m}$ is the maximum tolerable RTT of the ${m}$-th delay-sensitive task on the $({i},{j})$-th link. The indicator function $\mathds{1}(\mathcal{E})$ is $1$ if the event $\mathcal{E}$ is true, and 00 otherwise. Then, the average overall system cost is defined as the discounted summation of average system costs for all the scheduling periods, i.e.,
@@
Given the past scheduling actions and QoS observations (i.e., the system state), it is still difficult to accurately predict the relation between the scheduling action and task QoS in the current scheduling period. This is mainly because of the unknown interfering traffic and random channel contention. As a result, it is impossible to solve the above Bellman’s equations without any trial on the network performance. In this paper, we shall rely on the RL method to track the above unknown knowledge with the assistance of a preliminary observation dataset $\mathscr{S}^{s}$.
Particularly, before the optimization, the dataset $\mathscr{S}^{s}$ is collected from $M$ scheduling periods experiencing heterogeneous interfering traffic and link quality (e.g., the distances of links in $\mathcal{L}$ change due to mobility). Each of the scheduling periods (say the $\tau$-th one) is divided into two phases. In the first phase, a fixed testing scheduling action $\mathcal{A}^{p}$ is applied, and corresponding QoS observation $\mathcal{O}^{p}_{\tau}$ is obtained; in the second phase, a random scheduling action $\mathcal{A}^{s}_{\tau}$ according to certain distribution is applied, and another QoS observation $\mathcal{O}^{s}_{\tau}$ is obtained. Hence, the dataset $\mathscr{S}^{s}$ can be expressed as $\mathscr{S}^{s}\triangleq\left\{(\mathcal{O}^{p}_{\tau},\mathcal{A}^{p},\mathcal{O}^{s}_{\tau},\mathcal{A}^{s}_{\tau})|\tau=1,2,\ldots,M\right\}$.
@@ IV-A Performance Region Quantization
Note that the QoS observations of the testing scheduling action $\mathcal{A}^{p}$ should be collected to determine the performance region index of one scheduling period. In the online stage, one short slot can be reversed in each scheduling period to apply the testing scheduling action $\mathcal{A}^{p}$.
@@ V-A Offline Imitation Learning and Q-Network Training
$\mathcal{A}_{t}^{i}=\begin{cases}\operatorname*{arg\,min}UCB_{t}(k,\mathcal{A}^{i})&\text{with probability }1-\epsilon_{t},\\ \mathcal{A}^{i}\sim\mathrm{Unif}(\mathscr{A}^{i})&\text{with probability }\epsilon_{t},\end{cases}$ (18)

### v2

# ReinWiFi: Application-Layer QoS Optimization of WiFi Networks with Reinforcement Learning

@@ Abstract
The enhanced distributed channel access (EDCA) mechanism is used in current wireless fidelity (WiFi) networks to support priority requirements of heterogeneous applications. However, the EDCA mechanism can not adapt to particular quality-of-service (QoS) objective, network topology, and interference level. In this paper, a novel reinforcement-learning-based scheduling framework is proposed and implemented to optimize the application-layer quality-of-service (QoS) of a WiFi network with commercial adapters and unknown interference. Particularly, application-layer tasks of file delivery and delay-sensitive communication are jointly scheduled by adjusting the contention window sizes and application-layer throughput limitation, such that the throughput of the former and the round trip time of the latter can be optimized. Due to the unknown interference and vendor-dependent implementation of the WiFi adapters, the relation between the scheduling policy and the system QoS is unknown. Hence, a reinforcement learning method is proposed, in which a novel Q-network is trained to map from the historical scheduling parameters and QoS observations to the current scheduling action. It is demonstrated on a testbed that the proposed framework can achieve a significantly better performance than the EDCA mechanism.
@@ I Introduction
Reinforcement learning (RL) for radio resource management has been attracting tremendous attention since it is a promising technique to tackle unknown system statistics and solve the prohibitive policy optimization problem with tolerable complexity and good performance. Moreover, the RL technique also has great potential to optimize a wireless system even without accurate or complete observation of the system state, which might happen in practical implementations. Particularly, the optimization of WiFi systems with implementation constraints would be investigated in this paper.
@@
There have been a significant amount of works optimizing the throughput, delay or age-of-information (AoI) performance of wireless networks via the method of RL. Most of these works assumed full knowledge of the system state in algorithm design. They could be applied to the systems, where the global system state could be collected at a centralized controller in time. On the other hand, RL was also utilized to optimize the performance of wireless systems with distributive transmission scheduling, e.g., wireless fidelity (WiFi) systems. For instance, an adaptive channel contention mechanism was proposed for WiFi systems in [1], where a local RL agent was deployed at each user equipment (UE). The local agents adjusted the minimum contention window (MCW) size according to the global statistics of successful channel contention such that the transmission fairness among the agents can be ensured. In order to resolve the collision issue of distributive channel access, deep RL algorithms were proposed in [2] to determine the timing of doubling the contention window based on the estimated collision probability. In addition to the adaptive channel contention, a double deep Q-network (DDQN) [3] based rate adaptation algorithm was proposed in [4] to improve network throughput, where the agent inferred the optimal transmission rate based on the modulation and coding scheme (MCS) and frame loss rate. Most of the above literature assumed knowledge of the physical (PHY) layer and media access control (MAC) layer states. In fact, it might be challenging to obtain such knowledge in the scheduler design of a practical WiFi network. Moreover, the absence of knowledge on co-channel interference and the vendor-dependent implementation of WiFi adapters would also raise challenges. In [5], the rate of WiFi direct transmission was optimized by managing the transmission and receiving buffer in user-space, instead of the MAC layer buffer. However, the scheduling of WiFi transmission with implementation constraints has not been addressed satisfactorily.
@@
In this paper, we would like to shed some light on the RL-based scheduling design for practical WiFi systems suffering from unknown co-channel interference. Particularly, a framework, namely ReinWiFi, is proposed for the scheduling of delay-sensitive tasks and file delivery tasks in the application layer. In ReinWiFi, a controller collects the QoS observations of all the tasks periodically and determines the rate limitation and contention window size for all the transmitters such that the channel contention among them can be coordinated according to the overall QoS objective. Since there is no analytical model feasible for the above scheduling design, a novel Q-network is designed to make the scheduling decision. In order to accelerate the Q-learning, imitators are first proposed and trained to mimic the relationship between the scheduling action and QoS in different communication scenarios, respectively. Hence the Q-network can be trained with the imitators in an offline manner. It is shown by the experiments that the proposed framework can adapt to the variation of task number, interfering traffic, and link quality, and significantly outperforms the current EDCA mechanism defined in IEEE 802.11e.
@@ II-A Deployment Scenario
Because of the transmission latency constraint, the delay-sensitive tasks should be scheduled with higher priority than the file delivery ones. Hence, all the transmitters access the channel via the enhanced distributed channel access (EDCA) mechanism defined in IEEE 802.11e. Particularly, four access category (AC) queues, namely voice (VI), video (VO), best effort (BE), and background (BK), are adopted at all the transmitters. The transmission priorities of the four AC queues are differentiated by values of arbitration inter-frame spacing (AIFS) and contention window (CW) size. As in the practical systems, the file delivery tasks are scheduled with the BE priority, and the delay-sensitive tasks are scheduled with the VI priority. The latter has smaller AIFS and CW size, leading to a larger successful probability in channel contention. As a remark, due to the distributive channel contention mechanism, it is infeasible to accurately control the packet transmission order among the devices of a WiFi network with commercial WiFi adapter. Instead, the packet transmission in the ReinWiFi system is scheduled in a stochastic manner by adapting the CW sizes of AC queues in each device.
@@ II-B Task Queuing Model
For each file delivery task, all the information bits to be delivered are saved in an application-layer buffer, and a user datagram protocol (UDP) socket is established at the very beginning of transmission. The data dispatch from the buffer to the UDP socket is controlled by a dispatcher. The UDP socket encapsulates the received data from the dispatcher into UDP datagrams and forwards them to the driver of WiFi adapter.
As a remark, the new datagrams at the WiFi adapter may not be transmitted immediately. In fact, each WiFi adapter maintains four MAC-layer AC queues associated with the four transmission priorities, respectively. The arrival datagrams are saved in the corresponding queues and transmitted following unknown vendor’s protocol. The queuing status of the WiFi adapter is usually not accessible in the application-layer. Thus, it is infeasible for the proposed system to know when the WiFi adapter completely delivers a datagram; it is, therefore, infeasible for the proposed system to precisely control the transmission of a UDP datagram or an application-layer packet. As a result, the scheduling of the proposed system is designed based on the average observable performance in the application layer.
Specifically, the transmission time is organized into a sequence of scheduling periods, each with a duration of $T_{s}$ seconds. $T_{s}$ is sufficiently large to accommodate a number of MAC protocol data unit transmissions. Due to the invisibility of adapter status, the QoS of a file delivery task is measured by its application-layer throughput in one scheduling period. Particularly, for the ${m}$-th file delivery task of the $({i},{j})$-th link, its QoS in the $t$-th scheduling period $r_{{i},{j}}^{m}(t)$ is defined as the number of information bits transferred from the task buffer to the associated UDP socket. The dispatcher is designed to adaptively limit the throughput of the file delivery task such that delay-sensitive tasks could have a larger chance to access the channel. Hence, let $b_{{i},{j}}^{m}(t)$ be the throughput limitation of the ${m}$-th file delivery task of $({i},{j})$-th link in the $t$-th scheduling period, the dispatcher would make sure
@@
For each delay-sensitive task, a task queue and UDP socket are established at the very beginning. The application-layer packets arrive at the task queue periodically with a fixed average data rate. The first packet in the queue is forwarded to the UDP socket for WiFi transmission as long as the socket is idle. Due to the lack of MAC-layer status, the measurement of the transmission latency of a packet could hardly be accurate. Hence, we use the round-trip time (RTT) as the QoS measurement of delay-sensitive tasks. Particularly, for each delay-sensitive task, an acknowledgment will be sent back from the receiver to the transmitter when an application-layer packet is completely received. Hence, the transmitter can calculate the RTTs of all packet transmissions. For the ${m}$-th delay-sensitive task of the $({i},{j})$-th link ($\forall({i},{j})\in{\mathcal{L}},{m}\in{\mathcal{T}_{{i},{j}}^{r}},$), its QoS in the $t$-th scheduling period $d_{{i},{j}}^{m}(t)$ is defined as the average RTT of the packets transmitted in this scheduling period.
@@ II-C Scheduling Model
Denote the CW sizes of the VI and BE priorities of the ${i}$-th device in $t$-th scheduling period as $w_{i}^{\mathtt{VI}}(t)$ and $w_{i}^{\mathtt{BE}}(t)$ respectively, we focus on the joint scheduling of these channel contention parameters and the dispatchers’ throughput limitation $\{b_{{i},{j}}^{m}(t)|\forall({i},{j})\in{\mathcal{L}},{m}\in{\mathcal{T}_{{i},{j}}^{f}}\}$ in each scheduling period.
Particularly, each transmitter collects the QoS observations of its tasks in the end of each scheduling period and delivers them to a centralized controller, which can be implemented in an AP or other device, for making scheduling decision. Not all the tasks in the universal task sets are in the active state. The average RTTs and throughputs of inactive delay-sensitive and file delivery tasks are denoted by a sufficiently large value and 00, respectively. Hence, the aggregation of QoS observations received at the controller in the end of the $t$-th scheduling period can be represented as
@@
Due to the time-varying traffic of the interfering devices, the scheduling parameters, including the file throughput limitations and CW sizes, are adapted at the centralized controller in each schedule according to the system’s scheduling parameters and QoS observations in the past $N$ scheduling periods. Specifically, the aggregation of scheduling parameters in the $t$-th scheduling period ($\forall t$) is represented as
@@
Thus, in the very beginning of the $t$-th scheduling period, $\mathcal{A}_{t}$ ($\forall t$) is determined based on past scheduling parameters and QoS observations $\{(\mathcal{O}_{t-N},\mathcal{A}_{t-N}),(\mathcal{O}_{t-N+1},\mathcal{A}_{t-N+1}),\ldots,(\mathcal{O}_{t-1},\mathcal{A}_{t-1})\}$.
@@ III Problem Formulation
The proposed ReinWiFi system should successively make scheduling decisions for each scheduling period. Hence, it could be formulated as a Markov decision process (MDP).
@@
Denote $\mathcal{A}_{t}$ in (3) as the action in the $t$-th scheduling period, $\mathcal{A}_{t}^{i}\triangleq\left\{b_{i},{j}^{m}(t)|\forall j\in\mathcal{L}_{i},{m}\in{\mathcal{T}_{{i},{j}}^{f}}\right\}\cup\left\{w_{i}^{\mathtt{VI}}(t),w_{i}^{\mathtt{BE}}(t)\right\}$, as the local action of the $i$-th device in the $t$-th scheduling period. The scheduling policy $\Omega$ is a mapping from state space to action space as $\Omega(\mathcal{S}_{t})=\mathcal{A}_{t}$.
@@
where $\omega$ is a weight, $\mathrm{D}_{{i},{j}}^{m}$ is the maximum tolerable RTT of the ${m}$-th delay-sensitive task on the $({i},{j})$-th link. The indicator function $\mathds{1}(\mathcal{E})$ is $1$ if the event $\mathcal{E}$ is true, and 00 otherwise. Then, the overall system cost is defined as the average discounted summation of system costs for all the scheduling periods, i.e.,
@@
Given the system state, it is still difficult to accurately predict the relation between the scheduling action and task QoS in the current scheduling period. This is because of the unknown interfering traffic and random channel contention. It is therefore impossible to solve the above Bellman’s equations without any trial on the network performance. We shall rely on the RL method to track the above unknown knowledge with the assistance of a preliminary observation dataset $\mathscr{S}^{s}$.
Particularly, before the optimization, the dataset $\mathscr{S}^{s}$ is collected from $M$ scheduling periods experiencing heterogeneous interfering traffic and link quality. Each of the scheduling periods (say the $\tau$-th one) is divided into two phases. In the first phase, a fixed testing scheduling action $\mathcal{A}^{p}$ is applied, and corresponding QoS observation $\mathcal{O}^{p}_{\tau}$ is obtained; in the second phase, a random scheduling action $\mathcal{A}^{s}_{\tau}$ is applied, and another QoS observation $\mathcal{O}^{s}_{\tau}$ is obtained. Hence, the dataset $\mathscr{S}^{s}$ can be expressed as $\mathscr{S}^{s}\triangleq\left\{(\mathcal{O}^{p}_{\tau},\mathcal{A}^{p},\mathcal{O}^{s}_{\tau},\mathcal{A}^{s}_{\tau})|\tau=1,2,\ldots,M\right\}$.
@@ IV-A Performance Region Quantization
Note that the QoS observations of the testing scheduling action $\mathcal{A}^{p}$ should be collected to determine the performance region index of one scheduling period. In the online stage, one short period can be reversed in each scheduling period to apply the testing scheduling action $\mathcal{A}^{p}$.
@@ V-A Offline Imitation Learning and Q-Network Training
$\mathcal{A}_{t}^{i}=\begin{cases}\operatorname*{arg\,min}UCB_{t}(k,\mathcal{A}^{i})&\text{w.p. }1-\epsilon_{t},\\ \mathcal{A}^{i}\sim\mathrm{Unif}(\mathscr{A}^{i})&\text{w.p. }\epsilon_{t},\end{cases}$ (18)